Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours.

In the most recent month, 185,000 items were shipped to customers using 8,000 direct labor-hours. The company incurred a total of $27,600 in variable overhead costs.

According to the company’s standards, 0.04 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.50 per direct labor-hour.

Required:

1. What is the standard labor-hours allowed (SH) to ship 185,000 items to customers?

2. What is the standard variable overhead cost allowed (SH × SR) to ship 185,000 items to customers?

3. What is the variable overhead spending variance?

4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1
1 Standard labor-hours allowed 7400 =185000*0.04
2 Standard variable overhead cost allowed 25900 =7400*3.50
3 Variable overhead spending variance 1700 U =27600-25900
4 Variable overhead rate variance 400 F =27600-(8000*3.50)
Variable overhead efficiency variance 2100 U =3.50*(8000-7400)